OpenStudy (anonymous):

i have been stuck on this question for a week please help From 1985 to 2000 the percent p of total music sales with a certain format is modeled by p=5.7t - 11,308 where t is the year. (a) evaluate p for t= 1985 and t =2000 (b) use your results from part (a) to graph the equation from 1985 to 2000 (c) in what year was p = 74.9%? (a) p = for t = 1985 round to the nearest tenth as needed P = for t = 2000 round to the nearest tenth as needed (C) in what year was p = 74.9%?

OpenStudy (jdoe0001):

can you part A)?

OpenStudy (jdoe0001):

can you do that is, part A)

OpenStudy (anonymous):

huh

OpenStudy (jdoe0001):

(a) evaluate p for t= 1985 and t =2000 <---

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i don't understand

OpenStudy (jdoe0001):

"From 1985 to 2000 the percent p of total music sales with a certain format is modeled by \(\bf \large p=5.7t - 11,308\)" the equation to get the % of music sales per year is \(\bf \large p=5.7t - 11,308\) (a) evaluate p for t= 1985 and t =2000 that means "evaluate p" ,that is find the music sales in % "for t = 1985, and t = 2000" for the years of 1985 and 2000 how? well, we know the years in the equation is "t" so we set t = 1985 \(\bf \large p=5.7(1985) - 11,308\) t = 2000 \(\bf \large p=5.7(2000) - 11,308\)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

what is next?

OpenStudy (jdoe0001):

well, what "p" values did you get?

OpenStudy (anonymous):

are you suppose to subtract 1985 - 11,308

OpenStudy (jdoe0001):

\(\bf \bf \large p=5.7(1985) - 11,308 \implies p=5.7\times(1985) - 11,308 \)

OpenStudy (anonymous):

can you please explain i am not that good at math i need to know step by step

OpenStudy (jdoe0001):

not sure there many steps there, is just plain multiplication and addtion

OpenStudy (jdoe0001):

can't really kill a dead cat heheh, sorta speak